OpenStudy (anonymous):

Jenna buys 12 lb of meat. Some of the meat is chicken and costs $1.90/lb. The rest is beef, which costs $4.30/lb. She spends a total of $34.80. Which equation models this situation? Let c represent the number of pounds of chicken she buys. A. (1.9 + 4.3)(12 – c) = 34.8 B. 4.3c + 1.9(c + 14) = 34.8 C. 4.3c + 1.9(12) = 34.8 D. 1.9c + 4.3(12 − c) = 34.8

OpenStudy (loser66):

Any idea?

OpenStudy (anonymous):

D. or A.

OpenStudy (jdoe0001):

hint: \(\large \begin{array}{rllll} \textit{Jenna buys 12 lb of meat}\implies &b+c = 12 \\ \quad \\ \begin{array}{llll} \textit{chicken is 1.90, and beef is 4.30}\\ \textit{She spends a total of 34.80} \end{array}\implies &1.90c+4.30b=34.80 \end{array}\)

OpenStudy (jdoe0001):

use substitution from the 1st equation, and plug it in the 2nd equation

OpenStudy (anonymous):

So it is A. ?